OpenStudy (anonymous):

24a – 8 – 10a = -2(6 – 7a)

OpenStudy (anonymous):

i added the like terms

OpenStudy (anonymous):

but answer is crazy

OpenStudy (anonymous):

Combine like terms on the left side. You have 24a and you subtract 10a, you're left with 14a. On the right side distribute the -2. You now have: \[14a - 8 = -12 +14a\]

OpenStudy (anonymous):

i got 4.75

OpenStudy (anonymous):

24a – 8 – 10a = -2(6 – 7a) 14a =-12+14a+8 14a-14a=-4 0=-4? its weird -'_'-

OpenStudy (anonymous):

After you add 12 to both sides you have: \[4a+4=4a\] Divide both sides by 4: \[a+1=a\] Since no value is equal to itself plus one, there are no solutions.